SIR Q.PLEASE CAN YOU GIVE ME THE DERIVATIVE OF 1. ROOT SINx 2. (ax + b) to the power n.

SIR
Q.PLEASE CAN YOU GIVE ME THE DERIVATIVE OF
1. ROOT SINx
2. (ax + b) to the power n.

Grade:12th pass

1 Answers

vikas askiitian expert
13 years ago

1) y = (sinx)1/2

   dy/dx = d/dx[(sinx)1/2]

            = (1/2)(sinx)1/2-1 [d/dx(sinx)]                             ( d/dx(xn) = nxn-1)

            =(1/2)(sinx)-1/2(cosx)                                         (d/dx(sinx) =cosx )

2) y = (ax+b)n

 dy/dx = d/dx(ax+b)n

          =n(ax+b)n-1d/dx(ax+b)

         =n(ax+b)n-1[d/dx(ax) + d/dx(b)]                            (d/dx(constant) = 0)

        =n(ax+b)n-1 [a+0]

        na(ax+b)n-1